OpenStudy (anonymous):

solve polynomios subtract please 7.8x-3.4y+z -9.2x+4.8y-2.1z

OpenStudy (callisto):

Is it (i) (7.8x-3.4y+z) - (-9.2x+4.8y-2.1z) or (ii) (-9.2x+4.8y-2.1z) - (7.8x-3.4y+z ) ?

OpenStudy (anonymous):

it's (7.8x-3.4y+z) - (-9.2x+4.8y-2.1z)

OpenStudy (callisto):

(7.8x-3.4y+z) - (-9.2x+4.8y-2.1z) = (7.8x-3.4y+z) - (-9.2x) - (+4.8y) - (-2.1z) Can you simplify the brackets first?

OpenStudy (anonymous):

there is no brackets

OpenStudy (callisto):

Oh.. I mean those (..... ) - (....) For example: - (+4.8y) = -4.8y

OpenStudy (anonymous):

16.38x - 7.14y + 2.1z -9.2x + 4.8y - 2.1z ... correct me if I'm wrong

OpenStudy (callisto):

(7.8x-3.4y+z) - (-9.2x+4.8y-2.1z) = (7.8x-3.4y+z) - (-9.2x) -(+4.8y) -(-2.1z) = (7.8x-3.4y+z) + 9.2x - 4.8y + 2.1z = ... - (-9.2x) = (-1) (-9.2x) Negative times negative is positive, so you get = 9.2x Same for -(-2.1z) -(+4.8y) = (-1)(+4.8y) Negative times positive is negative, so you get = -4.8y Clear?

OpenStudy (anonymous):

clear

OpenStudy (callisto):

Now, group the like terms together, do the addition/subtraction for 7.8x-3.4y+z + 9.2x - 4.8y + 2.1z

OpenStudy (anonymous):

17.0x-8.2y+3.1z ...again correct me if I'm wrong

OpenStudy (callisto):

It's right :)

OpenStudy (anonymous):

another?

OpenStudy (anonymous):

The volume (v) of a sphere varies directly as the cube of its diameter (d). Write this statement in algebraic language, using an equation with the variables c, v, and d.

OpenStudy (callisto):

Start a new post for a new question!

OpenStudy (anonymous):

thank you fro you time:)